OpenStudy (anonymous):

Determine which of the following statements is true concerning the values described in statement # 1 and statement # 2. #1: The x-coordinate of the vertex of the equation y = 2x2 − 4x + 12 #2: The x-coordinate of the vertex of the equation y = 4x2 + 8x + 3 The value found in column #1 is greater than the value found in column #2. The value found in column #1 is less than the value found in column #2. The value found in column #1 is equivalent to the value found in column #2. The relationship between column #1 and column #2 cannot be determined by the information given.

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

do you know how to find the line of symmetry for a parabola...?

OpenStudy (campbell_st):

the reason is that the vertex lies on the line of symmetry.... and is the x value in the ordered pair

OpenStudy (anonymous):

this is the pre-test it's testing knowledge

OpenStudy (campbell_st):

yep.... but the question I'm asking is in a parabola \[y = ax^2 +bx + c\] do you know how to find the line of symmetry. the reason is that the vertex in on the line of symmetry and the value of x in the line of symmetry equation is the x value in the vertex... so its something that has some importance...

OpenStudy (anonymous):

I don't know because I haven't began learning about it.

OpenStudy (campbell_st):

lol... ok for a parabola \[y = ax^2 + bx + c\] the equation for the line of symmetry is \[x = \frac{-b}{2a}\] this will also give the value of the x value in the vertex so in your 1st question b= -4 and a = 2 2nd question b = 8, a = 4 so substitute them to get the corresponding x values... and then look at the statements and see which is true

OpenStudy (anonymous):

Ok thank you lots!! :D